OpenStudy (anonymous):

Medal and fan!! Alan can chop 3/4 of a carrot in 1/8 min. How many carrots can he chop in 1 min

OpenStudy (unanimoose):

Is it 3/8 or 3/4? The last question you opened said 3/8.

OpenStudy (anonymous):

3/4

OpenStudy (jhannybean):

Make a ratio of carrots chopped to minutes passed,

OpenStudy (unanimoose):

Ok, then just divide 3 and 4 and then multiply by 8.

OpenStudy (unanimoose):

\[\frac{ 3 }{ 4 } \times 8 = answer\]

OpenStudy (unanimoose):

But yeah he does chop pretty slow LOL.

OpenStudy (anonymous):

lol

OpenStudy (jhannybean):

\[\frac{3/4~\text{carrots}}{1/8 ~\text{mins}}=\frac{x~\text{carrots}}{1~\text{mins}}\]

OpenStudy (jhannybean):

\[\frac{3}{4}=\frac{1}{8}x\]\[x~\text{carrots}=\frac{3}{4}\cdot 8 = 2\cdot 3~\text{carrots}\]

OpenStudy (anonymous):

so its 6

OpenStudy (unanimoose):

3/4 = 0.75. 0.75 * 8 = carrots for the slowpoke.

OpenStudy (unanimoose):

Yes it's 6.